OpenStudy (e.mccormick):

Quick check on an even numbered problem: \[\mathrm{Given:}\;\vec{u}=(1,1,1),\, \vec{v}=(2,5,2), \langle \vec{u},\vec{v} \rangle=u_1v_1+2u_2v_2+u_3v_3\\ \langle \vec{u},\vec{v} \rangle=2+15+2\implies 19\\ \|\vec{u}\|=\sqrt{1+2+1}\implies 2\\ \|\vec{v}\|=\sqrt{4+50+4}\implies \sqrt{58}\\ d(\vec{u},\vec{v})=\sqrt{(2-1)^2+2(5-1)^2+(2-1)^2}\implies\sqrt{34}\]Correct?

OpenStudy (amistre64):

i take it that a dot product? the top looks like its more of a (uv)^2 setup

OpenStudy (amistre64):

but then that would be u.u + 2.u.v + v.v tho

OpenStudy (amistre64):

the length of (1,1,1) = sqrt(3), not 2 am i reading this right?

OpenStudy (e.mccormick):

The top is a given. Then I get to find \[\langle \vec{u},\vec{v} \rangle\\ \|\vec{u}\|\\ \|\vec{v}\|\\ d(\vec{u},\vec{v})\] As for that one, what about th2 2 in the given? Dosn't that change it to: \(\sqrt{1^2+2(1^2)+1^2}\)?

OpenStudy (amistre64):

its given that: \[\langle\vec{u},\vec{v} \rangle=u_1v_1+2u_2v_2+u_3v_3\\\] that is notation for dot product right?

OpenStudy (amistre64):

not sure why a 2 would be in there

OpenStudy (e.mccormick):

Ah, and I forgot to add it is an inner product defined in \(\mathbb{R}^n\). So they are defining the inner product as that.

OpenStudy (e.mccormick):

The 2 is there because it is given that way, as opposed to the Euclidian inner product which is just dot product.

OpenStudy (amistre64):

then wouldnt:\[(\vec u,\vec v)=1.2+2.1.5+1.2=2+10+2=14\]

OpenStudy (e.mccormick):

AH HA! Yes, I did make a mistake there! I don't know what I was thinking.

OpenStudy (amistre64):

and the norm is defined as: sqrt(u,v)

OpenStudy (amistre64):

..or something like that :)

OpenStudy (e.mccormick):

Yes, it is basically just the distance formula.

OpenStudy (e.mccormick):

But all it is measuring is the distance of a vector to itself, so you assume it is the distance to the origin, so the components of the vector are squared, added, then rooted.

OpenStudy (e.mccormick):

If you mean the \(d(\vec{u},\vec{v})\) part, that is distance between, not the norm.

OpenStudy (e.mccormick):

I think the rest of it is correct. I double checked the other numbers and the book to make sure I used the right formula. Thanks for finding the one I made a mistake on.

OpenStudy (amistre64):

youre welcome, ive had alot of practice finding all my own mistakes thru the years :)

OpenStudy (e.mccormick):

Oh, I know that feeling. You make a mistake and because you made it, it is invisible!